Problem 8.89 A turntable with a rotational inertia o.215 kg.m2 is rotating at 3.35 rad/s Suddenly, a disk with rotational inertia 0.106 kg m2 s dropped onto the turntable with its center on the rotation axis. Part A Assuming no outside forces act, whats the common rotational velocity of the turntable and disk? Express your answer with the appropriate units. w Value Units Submit My Answers Give Up
Answer #1

Rotational momentum is conserved.

initial = final = I*ωi = 0.215 * 3.35 kg.m^2/s = 0.72025 kg.m^2/s

final I = 0.215+0.106 = 0.32 kg.m^2

so final ω = 0.72025 / 0.32 rad/s = 2.25 rad/s